OpenStudy (anonymous):

You spent $4.30 on 3 yogurts and 2 juice boxes. A friend spent $2.60 on 1 juice box and 2 yogurts. What is the cost of 1 yogurt? What is the cost of 1 juice box? I just need help getting the problem set up, then I can do the rest by myself I think! :) Thank you.

OpenStudy (anonymous):

x = yogurts y = juice box 4.30 = 3x + 2y 2.60 = 2x + 1y always identify variables first.

OpenStudy (anonymous):

Can you tell me if I got it correct? Cost of 1 yogurt is .90 cents and the cost of a juice box is .80 cents.

OpenStudy (anonymous):

4.30 = .90(3) + 2(.80) 4.30 = 2.70 + 1.60 4.30 = 4.30 yes that works :)

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

you are welcome, good job!